### v1

# Hidden ferromagnetism of centrosymmetric antiferromagnets

Abstract
The time-reversal symmetry ($\mathcal{T}$) breaking is a signature of ferromagnetism, giving rise to such phenomena as the anomalous Hall effect (AHE) and orbital magnetism (OM). Nevertheless, $\mathcal{T}$ can be also broken in certain classes of antiferromagnets, such as weak ferromagnets or altermagnets, which remain invariant under the spatial inversion. In the light of this similarity with the ferromagnetism, it is tempting to ask whether such anomalous antiferromagnetic (AFM) state can be presented as a simplest ferromagnetic one, i.e. within a minimal unit cell containing only one magnetic cite. We show that such presentation is possible due to the special form of the spin-orbit (SO) interaction in an antiferroelectrically distorted lattice hosting this AFM state. The inversion symmetry, combined with the lattice translations, imposes a severe constraint on the form of the SO interaction, which becomes invariant under the symmetry operation $\{\mathcal{S}|{\bf t}\}$, combining the $180^{\circ}$ rotation of spins ($\mathcal{S}$) with the lattice shift $\bf t$}, connecting antiferromagnetically coupled sublattices. This is the fundamental symmetry property of centrosymmetric antiferromagnets, which justifies the use of the generalized Bloch theorem and transformation to the local coordinate frame with one magnetic cite per cell. It naturally explains the emergence of AHE and OM, and provides transparent expressions for these properties in terms of the electron hoppings and SO interaction operating between nearest neighbors as well as the orthorhombic strain of the next-nearest-neighbor hoppings. The idea is illustrated on a number of examples, using realistic models derived from first-principles calculations. These examples include two-dimensional square lattice, monoclinic VF4 and CuF2, and RuO2-type materials with the tetragonal symmetry.
I Introduction
The time-reversal symmetry ($\mathcal{T}$) breaking is a synonym of ferromagnetism. The magnetic unit cell of a regular ferromagnet coincides with the chemical one, but the electronic structure for the states with spins “up” and “down” is different, so that the ferromagnetic (FM) system is characterized by a finite spin magnetic moment, which has the same direction at all magnetic sites. On the contrary, the regular antiferromagnet is characterised by the doubling of the unit cell, so that two magnetic sublattices, which in the FM state would be connected by a primitive translation $\bf t$}, are occupied by atoms with opposite directions of spins. The basic symmetry of the regular antiferromagnetic (AFM) state is $\{\mathcal{T}|{\bf t}\}$, where $\mathcal{T}$ is combined with the lattice translation $\bf t$} of the chemical cell. Therefore, although $\mathcal{T}$ is microscopically broken, it is preserved macroscopically, after averaging over two AFM sublattices [1]. The electronic states of such antiferromagnets are Kramers degenerate.
Currently, a great deal of attention is paid to the systems, where the AFM alignment of spins coexists with some features, which are more characteristic for ferromagnets, including the macroscopic time-reversal symmetry breaking and splitting of AFM bands. Such systems are now called “altermagnets”, to emphasize the distinct character of these materials and their principal difference from the conventional ferromagnets and antiferromagnets [2, 3, 4].
On the other hand, the AFM materials with broken time-reversal symmetry have been known for decades. Particularly, in 1950s, using purely phenomenological symmetry arguments, Dzyaloshinskii pointed out that there are special types of antiferromagnets, which can host the phenomena of weak ferromagnetism [5], piezomagnetism [6], and magnetoelectricity [7]. Contrary to the regular antiferromagnetism, the magnetic unit cell of these material coincides with the chemical one, being the necessary precondition for macroscopic time-reversal symmetry breaking [1]. Therefore, it is clear that, since there cannot be two independent classes of AFM materials with broken $\mathcal{T}$, the phenomena considered by Dzyaloshinskii should be related to what is now called altermagnetism. In this respect, a very detailed phenomenological classification has been given by Turov, who suggested to divide unconventional antiferromagnets into two major classes, calling them “centrosymmetric” and “centroantisymmetric”, depending on whether the spatial inversion $\mathcal{I}$ enters the magnetic group alone or in the combination with $\mathcal{T}$ [8]. Thus, the centrosymmetric antiferromagnetism in Turov’s classification and encompassing such phenomena as weak ferromagnetism and piezomagnetism has a clear resemblance to what is now called altermagnetism, while the centroantisymmetric antiferromagnetism provides a general framework for understanding the magnetoelectricity. Besides the weak ferromagnetism, piezomagnetism, and magnetoelectricity, Turov consider a wide spectrum of phenomena expected in these unconventional antiferromagnets. Particularly, in 1962, Turov and Shavrov not only predicted the anomalous Hall effect (AHE) in centrosymmetric antiferromagnets, but explicitly stated that it can be driven directly by the AFM order [8, 9]. Turov also pointed out that since magnetoelectricity and weak ferromagnetism belong to two different classes, they are mutually exclusive and do not coexist unless they develop in different magnetic sublattices [10]. The typical example of such coexistence is GdFeO3, where Gd sublattice is magnetoelectric, while Fe sublattice is weakly ferromagnetic [11, 12].
Although many properties of altermagnetic materials were anticipated on the phenomenological level, the recent breakthrough in this field is related to the microscopic understanding of these properties, which became largely possible due to development of first-principles electronic structure calculations. The main attention is focused on the search of the new materials with the large splitting of AFM bands [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], which is regarded as a hallmark of altermagnetism [2, 3, 4, 23].
Such splitting is a consequence of the nonsymmorphic symmetry of crystals, which have several atoms in the chemical cell. The atomic positions in this case are generated by the symmetry operations $\{\mathcal{C}|{\bf t}\}$, combining the rotation $\mathcal{C}$ with the lattice shift $\bf t$}, which connects different sublattices. Then, the AFM order is obtained by combining some of the symmetry operations $\{\mathcal{C}|{\bf t}\}$ with $\mathcal{T}$, so that the magnetic unit cell remains the same as the chemical one, being in line with Dzyaloshinskii’s conjecture [1]. Nevertheless, the new aspect of the problem, which was overlooked in the earlier stages, is that the $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry gives rise to the spin-splitting of bands in the reciprocal space [2, 3, 23, 24]. This splitting is typically regarded as a signature of ferromagnetism in otherwise AFM materials and believed to be responsible for AHE and other phenomena, which are more typical for ferromagnets.
However, there is another important symmetry, $\mathcal{I}$, which was highlighted in Turov’s definition of “centrosymmetric antiferromagnetism”, but somewhat overshadowed by other symmetry properties in the modern developments of altermagnetism. Then, which symmetry is more important, for instance for the emergence of AHE in AFM substances: $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ or $\mathcal{I}$?
In our recent work, dealing with the minimal one-orbital model, where the spin-orbit (SO) interaction replicates the form of Dzyaloshinskii-Moriya (DM) interactions in the noncentrosymmetric bonds [5, 25], we have argues that the altermagnetic splitting of bands does not play a key role in AHE and orbital magnetism [26]. Namely, $\mathcal{T}$ can be broken even when the AFM bands are spin-degenerate. Similar behaviour has been found in a more sophisticated multi-orbital model for AFM $\kappa$-type organic conductors and orthorhombically distorted perovskites [17, 19]. Furthermore, we have argued that the fundamental symmetry of such AFM state is $\{\mathcal{S}|{\bf t}\}$, which combines the spin flip $\mathcal{S}=i\hat{\sigma}_{y}$ ($\hat{\sigma}_{y}$ being the Pauli matrix) with the lattice shift $\bf t$} connecting the antiferromagnetically coupled sublattices. This means that the eigenvectors for the spin states $\sigma=\pm$ differ only by a phase factor, which guarantees that (i) the spin bands are degenerate and (ii) the contributions of these bands to the anomalous Hall conductivity are equal to each other and, instead of the partial cancellation, which would occur in regular ferromagnets, we have an addition of such contributions [26].
The time-reversal operation itself is the combination of $\mathcal{S}$ and the complex conjugation $K$, $\mathcal{T}=\mathcal{S}K$. Therefore, if the microscopic Hamiltonian is complex, the operation $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$, which is expected for the regular antiferromagnets. That is why the spin degeneracy can exist even though the time-reversal symmetry is broken. Strictly speaking, the spin degeneracy in this case is not the Kramers’ degeneracy, because the latter implies that the system is invariant under $\mathcal{T}$, which is not the case here.
In this work we will explicitly show that the $\{\mathcal{S}|{\bf t}\}$ symmetry is directly related to the fact that $\mathcal{I}$ is conserved in the centrosymmetric antiferromagnets, which imposes a severe constraint on the form of the SO interaction in magnetic bonds. It is true that $\mathcal{I}$ transforms each sublattice to itself and does not interconnect the sites belonging to different sublattices [27]. However, from the viewpoint of magnetic bonds, $\mathcal{I}$, in the combination with the lattice translations, will unambiguously specify the properties of these bonds over the entire lattice: namely, knowing the bond properties around one magnetic site, one can find the properties around the neighboring site, belonging to different sublattice. Particularly, in the centrosymmetric antiferromagnets, the SO interactions around the sites of different sublattices differ by the sign. Thus, this SO interaction behaves like an AFM object, similar to the Néel field, which leads to the $\{\mathcal{S}|{\bf t}\}$ symmetry of microscopic Hamiltonian.
There is another important aspect of the $\{\mathcal{S}|{\bf t}\}$ symmetry: $\mathcal{S}$ is nothing but the $180^{\circ}$ rotation of spins and this rotation is combined with the shift $\bf t$}, which can be viewed as the lattice translation of a more compact unit cell containing one magnetic site. This constitutes the basis of the so-called generalized Bloch theorem [28], which states that by a unitary transformation to the local coordinate frame, where all spins are pointed in the positive direction of $z$, the AFM system can be described as it would have only one magnetic site in the unit cell, i.e. as a ferromagnet. The generalized Bloch theorem provides a mapping of the AFM system onto a FM one, which naturally explain the breaking of the time-reversal symmetry and all plethora of magnetic effect, which are originally known for the ferromagnets, but can be realized in some AFM systems.
The rest of the article is organized as follows. In Sec. II we will consider the basic symmetry properties of SO interaction imposed by $\mathcal{I}$ in the combination with lattice translations. Then, in Sec. III, we will show that the behavior of the SO interactions is related to even more fundamental properties of the magnetoelectric coupling in antiferroelectrically distorted lattice. Particularly, in order to preserve $\mathcal{I}$, the distortion must be antiferroelectric, which inevitably leads to the doubling of the unit cell and the sign change of the SO interaction when it is considered around two different sublattices. Then, in Sec. IV, we will apply the generalized Bloch theorem and show that, for some components of the SO interaction, the sign change can be compensated by the transformation to the local coordinate frame, so that the AFM system can be formally described as a FM one, with smaller unit cell containing only one magnetic site. In Sec. V, we will introduce the minimal model for the centrosymmetric antiferromagnets with broken $\mathcal{T}$ and derive transparent expressions for the anomalous Hall conductivity and orbital magnetization in the local coordinate frame. Applications for the square lattice, monoclinic VF4 and CuF2, as well as RuO2-type materials with the tetragonal symmetry will be considered in Sec. VI. Finally, in Sec. VII, we will summarize our work. Two appendices deal with the unitary transformation of the SO interaction and elimination of the same-sign components responsible for the weak spin ferromagnetism, and construction of the model Hamiltonian for VF4 and CuF2 on the basis of first-principles electronic structure calculations.
II Basic considerations
Let ${\bf T}_{1}$, ${\bf T}_{2}$, and ${\bf T}_{3}$ be the primitive translations of a crystal, consisting of two sublattices, which can be transformed to each other by a rotation, combined with the shift of the origin by ${\bf t}=\frac{1}{2}\left({\bf T}_{1}+{\bf T}_{2}+{\bf T}_{3}\right)$. Therefore, if one sublattice is centered at the origin ${\bf 0}=(0,0,0)$, another sublattice will be centered at $\bf t$}, and if $\bf 0$} is an inversion center, $\bf t$} is another inversion center [27]. Nevertheless, the midpoint between $\bf 0$} and $\bf t$} is not the inversion center, so that there is a finite DM interactions operating between different sublattices [5, 25].
The atomic positions in both the sublattices can be generated by the vectors ${\bf t}_{1}=\frac{1}{2}(-{\bf T}_{1}+{\bf T}_{2}+{\bf T}_{3})$, ${\bf t}_{2}=\frac{1}{2}({\bf T}_{1}-{\bf T}_{2}+{\bf T}_{3})$, and ${\bf t}_{3}=\frac{1}{2}({\bf T}_{1}+{\bf T}_{2}-{\bf T}_{3})$. However, ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ do not necessarily transform the DM interactions to themselves and in this sense are not the primitive translations. In any case, the atomic positions are fully specified by the vectors $\boldsymbol{R}=l{\bf t}_{1}+m{\bf t}_{2}+n{\bf t}_{3}$. If $l+m+n$ is even, the atom belong to the sublattice $1$. If it is odd, the atom belongs to the sublattice $2$. Thus, the DM interactions, $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, operating between different sublattices, are finite if $\boldsymbol{R}^{\prime}-\boldsymbol{R}$ is an odd superposition of ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$.
The DM interactions are induced by the SO coupling. In the simplest one-orbital model, the SO interactions are given by $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where $\hat{\boldsymbol{\sigma}}$ is the vector of Pauli matrices and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\sim\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ [29, 30]. Therefore, $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ obey the same symmetry principles.
[…]
$\boldsymbol{t}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{l^{\prime\prime}+m^{\prime\prime}+n^{\prime\prime}}\,\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}.$ (1)
Figure 1: Parameters of SO interaction around the atoms of two sublattices in the centrosymmetric structure obeying the $\{{\cal C}_{2x}|{\bf t}\}$ symmetry (a, top) and $\{{\cal C}_{2y}|{\bf t}\}$ symmetry (b, bottom). The atoms of the sublattices $1$ and $2$ are shown by filled and open circles, respectively. The bond directions, which are defined starting from the central site in the direction of neighboring sites, are shown by broken arrows. The vectors $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, attached to these bonds, are shown by the bold green arrows. According to this definition, the bonds around the sites 11 and $2$ have opposite directions, which also flip the directions of the vectors $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ attached to these bonds, as shown on the left and right parts of the figure. The unit cell is shown by the solid red line.
[…]
III Weak Ferromagnetism and Antiferroelectricity
In this section, we will elucidate even more fundamental reason why the DM interactions for the weak ferromagnets have a specific pattern shown in Fig. 1, which preserves the inversion symmetry, but doubles the unit cell, resulting in two sublattices. We believe is that the most fundamental quantity to consider is magnetoelectric coupling $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, relating the cross product of spins in the bond to the electric polarization in the same bond, $\vec{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{e}_{\boldsymbol{R}}$ is the unit vector in the direction of the spin moment at the site $\boldsymbol{R}$ [31, 32, 33].
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv[\mathcal{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}^{v,c}]$ is the $3\times 3$ tensor. According to our conventions [33], the bold character refers to the spin components ($c$), which couple to the cross product, while the vector symbol refers to the $x$, $y$, and $z$ components ($v$) of the electric polarization. Thus, under the spatial inversion, the $v$ components will transform according to the normal vector rules, while the $c$ components will transform according to the pseudovector rule.
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be finite even in centrosymmetric bonds. Indeed, the spatial inversion about the center of the bond yields the following property: ${\cal I}\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R}^{\prime},\boldsymbol{R}}=-\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, which means that $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ does not necessarily vanish. As the result, the electric polarization $\vec{P}$ can be induced by a noncollinear alignment of spins in otherwise centrosymmetric crystals [31, 32]. Alternatively, electric field $\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, acting in the bond, will interact with $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and makes the spins $\boldsymbol{e}_{\boldsymbol{R}}$ and $\boldsymbol{e}_{\boldsymbol{R}^{\prime}}$ noncollinear. The corresponding interaction energy is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the DM interaction induced by the electric field.
In the insulating state, $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be evaluated in terms of the superexchange theory, similar to the DM coupling. Indeed, the latter is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{2t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, while the former is $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\frac{2e}{V}\frac{t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, where e-e is the electron charge, $V$ is the unit cell volume, $U$ is the on-site Coulomb repulsion, $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is the transfer integral expanded in terms of the $2\times 2$ unity matrix $\hat{\mathbb{1}}$ and the vector of Pauli matrices, and $\hat{\vec{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{r}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is a similar expansion for the position operator [32].
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ obeys certain symmetry properties. To be specific, let us consider the ideal square lattice, which can be transformed to itself by the fourfold rotations around $z$ (${\cal C}_{4z}$) and the spatial inversion about the lattice sites. The primitive translations are ${\bf t}_{1}$ and ${\bf t}_{2}$, so that each lattice point is specified by $\boldsymbol{R}=m\,{\bf t}_{1}+n\,{\bf t}_{1}$. Furthermore, each bond obeys the twofold rotational symmetry about itself. Then, for the single bond along $x$, say 11-22 in Fig. 2, this tensor will have the following form [33]:
$\vec{\boldsymbol{\mathcal{P}}}_{1,2}=\left(\begin{array}[]{ccc}\mathcal{P}^{x,x}&0&0\\ 0&\mathcal{P}^{y,y}&\mathcal{P}^{y,z}\\ 0&\mathcal{P}^{z,x}&\mathcal{P}^{z,z}\end{array}\right).$
[…]
Figure 2: The form of magnetoelectric tensors $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the bonds around two neighboring sites in the ideal square lattice. The directions of the bonds are indicated by arrows.
The magnetoelectric tensor is periodic and, therefore, $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ for any $\boldsymbol{R}^{\prime\prime}$ (for instance, site $1$ in Fig. 2 can be moved to site $2$, etc.).
Next, we consider possible patters of electric fields caused by internal atomic displacements. For simplicity, we assume that all fields are parallel to $z$, but can have different directions in different bonds. For instance, such fields can be due to the buckling of the TM-L-TM bonds, where the intermediate ligand (L) sites are displaced parallel to $z$ relative to the transition-metal (TM) sites. Then, there are three main patterns, which are explained in Fig. 3:
  • (a)
[…]
  • (b)
[…]
  • (c)
    The antiferroelectric zigzag pattern, satisfying the property $\vec{E}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and doubling the unit cell as explained in Fig. 3(c). Nevertheless, these electric fields respect the inversion symmetry, so as the vectors of DM interactions. As far as the DM interactions is concerned, the magnetic texture is commensurate and characterized by a canting of spins. The typical example is the weak ferromagnetism [5], realized in La2CuO4 and other materials with orthorhombically distorted perovskite structure [48, 11].
Figure 3: Main patters of internal electric fields (left) and corresponding to them Dzyaloshinskii-Moriya interactions (right) on the square lattice: (a) ferroelectric, (b) antiferroelectric noncentrosymmetric, and (c) antiferroelectric centrosymmetric. The doubled unit cell in the latter case is shown by broken line. xyxy is the coordinate frame of the regular square lattice. $x^{\prime}y^{\prime}$ is the coordinate frame, which is typically used for the doubled unit cell, such as in Fig. 1.
To summarize this section, in the weak FM mode, the DM interactions respect the inversion symmetry. This means that the lattice is inevitably antiferroelectric, which results in the doubling of the unit cell. For two other modes, the DM interactions are translationally invariant, but the inversion symmetry is broken.
IV Generalized Bloch Theorem
Suppose that two sublattices are ordered antiferromagnetically. The corresponding AFM order can be described by the propagation vector $\boldsymbol{q}={\bf G}_{k}$, where ${\bf G}_{k}=\frac{2\pi}{V}\varepsilon_{ijk}[{\bf T}_{i}\times{\bf T}_{j}]$ is one of the reciprocal lattice translations ($k=$ 11, 22, or $3$) and $\varepsilon_{ijk}$ is the antisymmetric Levi-Civita symbol, so that $\boldsymbol{q}\cdot\boldsymbol{R}=0$ and $\pi$ ($\rm mod$} $2\pi$) for the sublattices $1$ and $2$, respectively.
Then, let us assume that the spins lie in the $xy$ plane, forming the angle $\alpha$ with the axis $x$. Corresponding Néel field is given by $\pm B(\cos\alpha\hat{\sigma}_{x}+\sin\alpha\hat{\sigma}_{y})$, where the signs ++ and - stand for the sublattices $1$ and $2$, respectively.
Consider the symmetry operations $\{\hat{U}_{\boldsymbol{R}}|\boldsymbol{R}\}$, combining the lattice shift $\boldsymbol{R}$ with the $\rm SU(2)$} rotation of spins
$\hat{U}_{\boldsymbol{R}}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{rr}1&-1\\ 1&1\end{array}\right)\left(\begin{array}[]{cc}e^{\frac{i(\boldsymbol{q}\cdot\boldsymbol{R}+\alpha)}{2}}&0\\ 0&e^{-\frac{i(\boldsymbol{q}\cdot\boldsymbol{R}+\alpha)}{2}}\end{array}\right)$ (2)
to the local coordinate frame, where the field is $B\hat{\sigma}_{z}$ at both magnetic sublattices. Thus, from the viewpoint of magnetic field in the local coordinate frame, the system behaves as a ferromagnet. The result is well known in the theory of spin-spiral magnetism [28, 49]. Our next goal is to find whether the same applies for the transfer integrals $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where the first and second terms are, respectively, even and odd in the SO coupling. To the lowest order, $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the regular transfer integral without the SO coupling, while $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is induced by the SO coupling (or simply the SO coupling). Therefore, we have to find how $\hat{\boldsymbol{\sigma}}$ and $\hat{\mathbb{1}}$ are transformed by $\hat{U}_{\boldsymbol{R}}$ and $\hat{U}_{\boldsymbol{R}^{\prime}}$. These transformations are given by:
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{x}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=-\cos\left(\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}+\alpha\right)\hat{\sigma}_{z}-\sin\left(\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}+\alpha\right)\hat{\sigma}_{y},$ (3)
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{y}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=-\sin\left(\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}+\alpha\right)\hat{\sigma}_{z}+\cos\left(\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}+\alpha\right)\hat{\sigma}_{y},$ (4)
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{z}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=i\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\mathbb{1}}+\cos\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\sigma}_{x},$ (5)
and
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=\cos\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\mathbb{1}}+i\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\sigma}_{x}.$ (6)
Considering the transfer integrals and the SO interaction between different sublattices, we have $\frac{\boldsymbol{q}\cdot(\boldsymbol{R}\pm\boldsymbol{R}^{\prime})}{2}=\frac{\pi}{2}$ ($\rm mod$} $\pi$) and, therefore,
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{x}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=-\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}\left(\cos\alpha\hat{\sigma}_{y}-\sin\alpha\hat{\sigma}_{z}\right),$ (7)
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{y}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=-\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}+\boldsymbol{R}^{\prime})}{2}\left(\sin\alpha\hat{\sigma}_{y}+\cos\alpha\hat{\sigma}_{z}\right),$ (8)
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{z}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=i\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\mathbb{1}}$, (9)
and
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=i\sin\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\sigma}_{x}.$ (10)
Thus, we have the following property: $\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}$ for any $\boldsymbol{R}^{\prime\prime}$. The same holds for $\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{z}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}$. On the other hand, for $a=$ $x$ and $y$ we have $\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\sigma}_{a}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=(-1)^{l^{\prime\prime}+m^{\prime\prime}+n^{\prime\prime}}\,\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{a}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}$, which is the same as the property of SO interaction parameters $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ given by Eq. (1).
Combining these transformations of $\hat{\mathbb{1}}$ and $\hat{\boldsymbol{\sigma}}$ with $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=(t^{x}_{\boldsymbol{R},\boldsymbol{R}^{\prime}},t^{y}_{\boldsymbol{R},\boldsymbol{R}^{\prime}},t^{z}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$, it is straightforward to see that
$t_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}\,\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\,\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger},$ (11)
$t^{x}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}\,\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\sigma}_{x}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=t^{x}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\,\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{x}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger},$ (12)
$t^{y}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}\,\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\sigma}_{y}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=t^{y}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\,\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{y}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger},$ (13)
and
$t^{z}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}\,\hat{U}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime}}^{\phantom{\dagger}}\hat{\sigma}_{z}\hat{U}_{\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}^{\dagger}=(-1)^{l^{\prime\prime}+m^{\prime\prime}+n^{\prime\prime}}t^{z}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\,\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\sigma}_{z}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}.$ (14)
Thus, in the local coordinate frame, not only the regular transfer integrals, but also two components of the SO interaction appear to be periodic on the lattice specified by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$, and transforming the sublattices $1$ and $2$ to each other. The third ($z$) component of the SO interactions is not periodic on this minimal lattice and would generally require to use the two-sublattice unit cell specified by ${\bf T}_{1}$, ${\bf T}_{2}$, and ${\bf T}_{3}$. However, for the weak ferromagnets, at least one component of the SO interaction should have the same sign in all the bonds around each magnetic site and, therefore, can be eliminated [29, 26, 50]. By properly specifying the quantization axis, this component can be always chosen as $z$. The elimination procedure is valid to the first order in the SO coupling and its details are explained in Appendix A. In some systems, like RuO2 and La2CuO4, having, respectively, $P4_{2}/mnm$ and $Bmab$ symmetry, the $z$ components of the SO coupling is equal to zero and, therefore, no elimination is required. For comparison, the sign-alternating components of the SO coupling do not contribute to the weak ferromagnetism and cannot be eliminated simultaneously in all the bonds [26].
For the transfer integrals operating within the sublattices, the transformation to the local coordinate frame is given by
$\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}\hat{\mathbb{1}}\hat{U}_{\boldsymbol{R}^{\prime}}^{\dagger}=\cos\frac{\boldsymbol{q}\cdot(\boldsymbol{R}-\boldsymbol{R}^{\prime})}{2}\hat{\mathbb{1}}$, (15)
which is the same for the sublattices $1$ and $2$. Therefore, these sublattices can be transformed to each other by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ if the transfer integrals $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ operating in the sublattice $1$ are the same as in the sublattice $2$. In this case, there will be no altermagnetic band splitting, originating from the difference of $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the sublattices $1$ and $2$ [26, 51]. However, this splitting does not play a key role in AHE, which emerges already in the spin-degenerate bands [26]. Therefore, as the first approximation, the change of the transfer integrals related to the altermagnetic band splitting can be neglected.
To summarize this section, the minimal model for AHE in weak ferromagnets can be formulated in the local coordinate frame, combining the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ with the $\rm SU(2)$} rotations of spins. In the weak ferromagnets, the transformation to the local coordinate frame compensates the effect of the antiferroelectric order, which resulted in the doubling of the unit cell (see Sec. III). As the result, the electronic Hamiltonian in this local coordinate frame becomes periodic on the lattice specified by ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$. Thus, the AFM system with the SO coupling can be effectively describe as a FM one with only one site in the unit cell.
V Minimal model
AHE in centrosymmetric antiferromagnets has been predicted several decades ago on phenomenological grounds [8, 9]. In 1997, the magneto-optical effect (the ac analog of AHE) in weak ferromagnets has been studied quantitatively on the basis of first-principles electronic structure calculations [52]. Particularly, it was shown that the effect is strong and comparable with the one in regular FM state (though the net spin magnetization in the weak and regular FM states differed by two orders of magnetide). This behavior was attributed to the form of orbital magnetization, which substantially deviates from the spin one. In this section, we introduce the minimal model, describing AHE and orbital magnetization on the microscopic level.
V.1 Hamiltonian
Using the generalized Bloch theorem, the minimal model for AHE in the reciprocal space can be formulated in the local coordinate frame via the Fourier transform, which combines the Bloch factor, $e^{i\boldsymbol{k}\cdot\boldsymbol{R}}$, associated with the translation $\boldsymbol{R}$, with additional prefactors arising from the $\rm SU(2)$} rotation of spins, $\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}$, with the propagation vector $\boldsymbol{q}={\bf G}_{k}$ [28, 49]. Suppose that $y$ component of the SO coupling is sign-alternating, while $x$ and $z$ components have the same sign and can be eliminated, as explained in Appendix A. As it will become clear below, this is rather general situation, relevant to the behavior of many weak ferromagnets. The corresponding $\times22$ Hamiltonian is given by:
$\hat{\cal H}_{\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}\hat{\mathbb{1}}+h^{1}_{\boldsymbol{k}}\hat{\sigma}_{x}-h^{y}_{\boldsymbol{k}}\hat{\sigma}_{y}+B\hat{\sigma}_{z},$ (16)
where $h^{1}_{\boldsymbol{k}}$ and $h^{y}_{\boldsymbol{k}}$ are the Fourier images of the nearest-neighbor hoppings and sign-alternating part of the SO coupling, respectively, and $h^{\phantom{0}}_{\boldsymbol{k}}$ is the Fourier image of transfer integrals within the sublattice, assuming that these transfer integrals have the same form in both AFM sublattices and the altermagnetic part, responsible for the spin splitting, can be neglected. These Fourier images can be obtained using Eqs. (10), (8) (for $\alpha=\pi/2$), and (15), respectively. The Néel field is also supposed to be parallel to $y$ ($\alpha=\pi/2$), the same as the sign-alternating part of the SO coupling. Alternatively, one can consider the situation where the sign-alternating part of the SO coupling and the Néel field are parallel to $x$ ($\alpha=0$), which yields the same Hamiltonian in the local coordinate frame – see Eqs. (7) and (8) for $\alpha=0$ and $\pi/2$, respectively.
The Hamiltonian has two eigenvalues:
$\varepsilon^{\pm}_{\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}\mp\sqrt{B^{2}+(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}\equiv h^{\phantom{0}}_{\boldsymbol{k}}\mp A_{\boldsymbol{k}},$ (17)
for the majority ($\varepsilon^{+}_{\boldsymbol{k}}$) and minority ($\varepsilon^{-}_{\boldsymbol{k}}$) spin states in the local coordinate frame. Searching corresponding to them eigenvectors in the form
$|u^{+}_{\boldsymbol{k}}\rangle=\left(\begin{array}[]{c}\cos\theta_{\boldsymbol{k}}e^{i\phi_{\boldsymbol{k}}}\\ \sin\theta_{\boldsymbol{k}}\end{array}\right)\equiv|u^{\phantom{+}}_{\boldsymbol{k}}\rangle$ (18)
and
$|u^{-}_{\boldsymbol{k}}\rangle=\left(\begin{array}[]{c}-\sin\theta_{\boldsymbol{k}}\\ \cos\theta_{\boldsymbol{k}}e^{-i\phi_{\boldsymbol{k}}}\end{array}\right),$ (19)
it is straightforward to find that
$A_{\boldsymbol{k}}\cos 2\theta_{\boldsymbol{k}}=B$ (20)
and
$A_{\boldsymbol{k}}\sin 2\theta_{\boldsymbol{k}}e^{i\phi_{\boldsymbol{k}}}=h^{1}_{\boldsymbol{k}}+ih^{y}_{\boldsymbol{k}},$ (21)
where
$\theta_{\boldsymbol{k}}=\frac{1}{2}\arctan\frac{\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}}{B}$ (22)
($0\leq\theta_{\boldsymbol{k}}<\pi$) and
$\phi_{\boldsymbol{k}}=\arctan\left(\frac{h^{y}_{\boldsymbol{k}}}{h^{1}_{\boldsymbol{k}}}\right)$ (23)
($0\leq\phi_{\boldsymbol{k}}<2\pi$).
V.2 Anomalous Hall Conductivity
Using the above expressions and considering the situation when $B$ is sufficiently large, so that the majority-spin band is partly occupied, while the minority-spin band is empty, it is straightforward to obtain that the expression for the Berry curvature, $\Omega^{c}(\boldsymbol{k})=-2{\rm Im}\left\langle\partial_{k_{a}}u_{\boldsymbol{k}}|\partial_{k_{b}}u_{\boldsymbol{k}}\right\rangle$, can be rearranged as
$\Omega_{\boldsymbol{k}}^{c}=\sin 2\theta_{\boldsymbol{k}}\left(\partial_{k_{a}}\theta_{\boldsymbol{k}}\partial_{k_{b}}\phi_{\boldsymbol{k}}-\partial_{k_{a}}\phi_{\boldsymbol{k}}\partial_{k_{b}}\theta_{\boldsymbol{k}}\right),$ (24)
where
$\sin 2\theta_{k}=-\frac{\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}}{A_{\boldsymbol{k}}},$ (25)
$\partial_{k_{a}}\theta_{\boldsymbol{k}}=\frac{1}{2}\frac{(h^{1}_{\boldsymbol{k}}\partial_{k_{a}}h^{1}_{\boldsymbol{k}}+h^{y}_{\boldsymbol{k}}\partial_{k_{a}}h^{y}_{\boldsymbol{k}})B}{A_{\boldsymbol{k}}^{2}\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}},$ (26)
$\partial_{k_{a}}\phi_{\boldsymbol{k}}=\frac{h^{1}_{\boldsymbol{k}}\partial_{k_{a}}h^{y}_{\boldsymbol{k}}-h^{y}_{\boldsymbol{k}}\partial_{k_{a}}h^{1}_{\boldsymbol{k}}}{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}},$ (27)
and abcabc is an even permutation of xyzxyz.
Then, after some algebra, $\Omega_{\boldsymbol{k}}^{c}$ can be further transformed to
$\Omega_{\boldsymbol{k}}^{c}=\frac{B}{2A^{3}_{\boldsymbol{k}}}\left(\partial_{k_{a}}h^{y}_{\boldsymbol{k}}\partial_{k_{b}}h^{1}_{\boldsymbol{k}}-\partial_{k_{a}}h^{1}_{\boldsymbol{k}}\partial_{k_{b}}h^{y}_{\boldsymbol{k}}\right).$ (28)
The anomalous Hall conductivity is given by the Brillouin zone ($\rm BZ$}) integral
$\sigma_{ab}=-\frac{1}{V}\int_{\rm BZ}\frac{d\boldsymbol{k}}{\Omega_{\rm BZ}}f_{\boldsymbol{k}}\Omega^{c}_{\boldsymbol{k}},$
where $\Omega_{\rm BZ}=\frac{(2\pi)^{3}}{V}$ is the $\rm BZ$} volume and $f_{\boldsymbol{k}}$ is the Fermi-Dirac distribution function for $\varepsilon^{+}_{\boldsymbol{k}}$.
V.3 Orbital Magnetization
The orbital magnetization along $c$ is given by the $\rm BZ$} integral of
${\cal M}^{c}_{\boldsymbol{k}}={\rm Im}\left\langle\partial_{k_{a}}u_{\boldsymbol{k}}|\hat{\cal H}_{\boldsymbol{k}}+\varepsilon^{+}_{\boldsymbol{k}}-2\varepsilon_{\rm F}|\partial_{k_{b}}u_{\boldsymbol{k}}\right\rangle$ (29)
with the Fermi-Dirac distribution function [26, 53, 54]. Using the explicit form of $\hat{\cal H}_{\boldsymbol{k}}$, $\varepsilon^{+}_{\boldsymbol{k}}$, and $|u_{\boldsymbol{k}}\rangle$, ${\cal M}^{c}_{\boldsymbol{k}}$ can be rearranged as the sum of the following four contributions:
${\cal M}_{\boldsymbol{k},{\rm I}}^{c}=-\frac{1}{2}\sin 2\theta_{\boldsymbol{k}}\left(2h^{\phantom{0}}_{\boldsymbol{k}}-A_{\boldsymbol{k}}\right)\left(\partial_{k_{a}}\theta_{\boldsymbol{k}}\partial_{k_{b}}\phi_{\boldsymbol{k}}-\partial_{k_{a}}\phi_{\boldsymbol{k}}\partial_{k_{b}}\theta_{\boldsymbol{k}}\right),$ (30)
${\cal M}_{\boldsymbol{k},{\rm II}}^{c}=\left(\varepsilon_{\rm F}-\frac{B}{2}\right){\Omega}_{\boldsymbol{k}}^{c},$ (31)
${\cal M}_{\boldsymbol{k},{\rm III}}^{c}=h^{1}_{\boldsymbol{k}}\cos^{2}\theta_{\boldsymbol{k}}\cos\phi_{\boldsymbol{k}}\left(\partial_{k_{a}}\theta_{\boldsymbol{k}}\partial_{k_{b}}\phi_{\boldsymbol{k}}-\partial_{k_{a}}\phi_{\boldsymbol{k}}\partial_{k_{b}}\theta_{\boldsymbol{k}}\right),$ (32)
and
${\cal M}_{\boldsymbol{k},{\rm IV}}^{c}=h^{y}_{\boldsymbol{k}}\cos^{2}\theta_{\boldsymbol{k}}\sin\phi_{\boldsymbol{k}}\left(\partial_{k_{a}}\theta_{\boldsymbol{k}}\partial_{k_{b}}\phi_{\boldsymbol{k}}-\partial_{k_{a}}\phi_{\boldsymbol{k}}\partial_{k_{b}}\theta_{\boldsymbol{k}}\right).$ (33)
Then, using Eq. (24), all these terms can be related to $\Omega_{\boldsymbol{k}}^{c}$ as
${\cal M}_{\boldsymbol{k},{\rm I}}^{c}=-\frac{1}{2}\left(2h^{\phantom{0}}_{\boldsymbol{k}}-A_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c},$ (34)
${\cal M}_{\boldsymbol{k},{\rm III}}^{c}=\frac{h^{1}_{\boldsymbol{k}}\cos\phi_{\boldsymbol{k}}\cos^{2}\theta_{\boldsymbol{k}}}{\sin 2\theta_{\boldsymbol{k}}}\Omega_{\boldsymbol{k}}^{c},$ (35)
and
${\cal M}_{\boldsymbol{k},{\rm IV}}^{c}=\frac{h^{y}_{\boldsymbol{k}}\sin\phi_{\boldsymbol{k}}\cos^{2}\theta_{\boldsymbol{k}}}{\sin 2\theta_{\boldsymbol{k}}}\Omega_{\boldsymbol{k}}^{c}.$ (36)
Noting that
$\frac{\cos^{2}\theta_{\boldsymbol{k}}}{\sin 2\theta_{\boldsymbol{k}}}=\frac{1}{2}\frac{B-A_{\boldsymbol{k}}}{\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}}$ (37)
and $h^{1}_{\boldsymbol{k}}\cos\phi_{\boldsymbol{k}}+h^{y}_{\boldsymbol{k}}\sin\phi_{\boldsymbol{k}}=\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}$, it is straightforward to see that ${\cal M}_{\boldsymbol{k},{\rm III}}^{c}+{\cal M}_{\boldsymbol{k},{\rm IV}}^{c}=\frac{1}{2}\left(B-A_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c}$. Summing up all four terms, one can get the final expression:
${\cal M}^{c}_{\boldsymbol{k}}=\left(\varepsilon_{\rm F}-h^{\phantom{0}}_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c}.$ (38)
Thus, the orbital magnetization is
${\cal M}^{c}=\int_{\rm BZ}\frac{d\boldsymbol{k}}{\Omega_{\rm BZ}}f_{\boldsymbol{k}}\left(\varepsilon_{\rm F}-h^{\phantom{0}}_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c}.$
VI Examples
VI.1 Square lattice of perovskites
As the first example, let us consider the single layer of orthorhombically distorted perovskites with the $Pbnm$ or $Bmab$ symmetry [26]. The lattice is specified by the primitive translations ${\bf T}_{1}=(1,0,0)$ and ${\bf T}_{2}=(0,1,0)$, and there are two sublattices, which are connected by the vectors ${\bf t}_{1}=(\frac{1}{2},\frac{1}{2})$ and ${\bf t}_{2}=(-\frac{1}{2},\frac{1}{2})$. The symmetry operation transforming two sublattices to each other is $\{{\cal C}_{2x}|{\bf t}_{1}\}$, meaning that $x$ components of the SO coupling is sign-alternating. For the space group $Bmab$, the lattice is additionally invariant under the mirror reflection $y\to-y$.
The minimal model includes the following ingredients [26]: (i) The hoppings between first, second, and third nearest neighbors ($t_{1}$, $t_{2}$, and $t_{3}$, respectively); (ii) The orthorhombic strain of the second-nearest neighbor hoppings, $\delta t_{2}$, such that the total hopping along $x$ and $y$ is, respectively, $t_{2}+\delta t_{2}$ and $t_{2}-\delta t_{2}$ in both sublattices; (iii) The SO coupling in the noncentrosymmetric nearest-neighbor bonds, which can be taken in the form $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\pm it_{x}\hat{\sigma}_{x}$. The $y$ and $z$ components of $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, responsible for the weak spin ferromagnetism, can be eliminated via the unitary transformation [29, 50], as explained in Appendix A. We would like to reiterate again that the SO interaction obeys the following property, being the consequence of the inversion symmetry and translational invariance: $\hat{\cal H}^{\rm so}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\,\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. The regular SO interaction is not periodic on the lattice, specified by ${\bf t}_{1}$ and ${\bf t}_{2}$. However, it can be made periodic in the local coordinate frame as explained in Secs. II-IV; (iv) The Néel field, which is also parallel to $x$, $(-1)^{m+n+1}\,B\hat{\sigma}_{x}$, where odd and even values of mm++$n$ correspond to the sublattices $1$ and $2$, respectively. This term is transformed to $B\hat{\sigma}_{z}$ in the local coordinate frame, which is specified by the transformation (2) with $\alpha=0$ and $\boldsymbol{q}=(2\pi,0)$. We do not consider the altermagnetic deformation of the third-neighbor hoppings, which is identically equal to zero for the $Bmab$ symmetry of La2CuO4 and expected to be small for other materials [26].
Thus, one can readily find the Hamiltonian (16) in the local coordinate frame, where $h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{\boldsymbol{k}}$, and $h^{y}_{\boldsymbol{k}}$ are obtained by combining the Fourier transforms with the matrix elements of $\hat{\mathbb{1}}$ and $\hat{\sigma}_{x}$ given by Eqs. (10) and (7), respectively. Moreover, it is convenient to shift the $\boldsymbol{k}$-mesh: $(k_{x},k_{y})\to(k_{x}+\pi,k_{y})$. Altogether, this yields: $h^{\phantom{0}}_{\boldsymbol{k}}=2t_{2}(\cos k_{x}+\cos k_{y})+2\delta t_{2}(\cos k_{x}-\cos k_{y})+4t_{3}\cos k_{x}\cos k_{y}$, $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, and $h^{y}_{\boldsymbol{k}}=-4t_{x}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$. One can clearly see that, after replacing the Pauli matrices $\boldsymbol{\sigma}$ by pseudospin matrices $\boldsymbol{\tau}$, the obtained Hamiltonian is the same as the one in global coordinate frame for the $\sigma=$- states and $\delta t_{3}=0$, which is given by Eq. (1) of Ref. [26]. Obviously, it will yield the same anomalous Hall conductivity and orbital magnetization. Nevertheless, the Berry curvature can be now written in the more compact form, using Eq. (28):
$\Omega_{\boldsymbol{k}}^{z}=\frac{Bt_{1}t_{x}}{A_{\boldsymbol{k}}^{3}}\left(\cos k_{x}-\cos k_{y}\right).$ (39)
The results are summarized in Fig. 4.
Figure 4: Results for the square lattice model with the parameters (unless it is specified otherwise) $t_{1}=-1$, $t_{2}=-t_{3}=0.1$, and $\delta t_{2}=-t_{x}=0.05$ (all are in units of $B$): (a) Berry curvature, $\Omega^{z}_{\boldsymbol{k}}$, in the first Brillouin zone; (b) Similar plot for orbital magnetization, ${\cal M}^{z}_{\boldsymbol{k}}$, corresponding to $n_{\rm el}=0.5$ electrons; Band filling dependence of (c) the anomalous Hall conductivity, $\sigma_{xy}$, and (d) the orbital magnetization, ${\cal M}^{z}$; (e) Fermi surface at $n_{\rm el}=0.5$ with (red) and without (blue) the orthorhombic strain $\delta t_{2}$; and (f) The orthorhombic strain dependence of $\sigma_{xy}$ and ${\cal M}^{z}$ for $n_{\rm el}=0.9$.
In comparison with calculations in the global coordinate frame [26], the generalized Bloch theorem allows us to use smaller unit cell. Therefore, the first Brillouin zone becomes larger, where $\Omega^{z}_{\boldsymbol{k}}$ has two positive and two negative areas [Fig. 4(a)], which tend to compensate each other. Therefore, in order to obtain finite $\sigma_{xy}$ and ${\cal M}^{z}$, it is essential to consider the orthorhombic strain, $\delta t_{2}$ [17, 26]. This strain deforms the Fermi surface, as shown in Fig. 4(e), so that the positive and negative areas of $\Omega^{z}_{\boldsymbol{k}}$ will contribute to $\sigma_{xy}$ with different weights, resulting in finite $\sigma_{xy}$. According to Eq. (38), in the case of ${\cal M}^{z}$, there is an additional modulation of $\Omega^{z}_{\boldsymbol{k}}$ by $h^{\phantom{0}}_{\boldsymbol{k}}$, which makes inequivalent the positive and negative areas of ${\cal M}^{z}_{\boldsymbol{k}}$ [see Fig. 4(b)]. Nevertheless is small if $\delta t_{2}$ is small.
Finally, AHE and orbital magnetization can be efficiently controlled by the orthorhombic strain, where the transition from tensile ($\delta t_{2}<0$) to compressive ($\delta t_{2}>0$) strain changes the sign of $\sigma_{xy}$ and ${\cal M}^{z}$ [see Fig. 4(f)]. This effect has a clear similarity with the piezomagnetizm [18, 22, 26]. For small $\delta t_{2}$, the contribution $\varepsilon_{\rm F}\Omega^{z}_{\boldsymbol{k}}$ to the orbital magnetization dominates and ${\cal M}^{z}\sim\sigma_{xy}$.
VI.2 VF4 and CuF2
As the next example, we consider the minimal model for VF4 and CuF2, which are regarded as altermagnetic candidates [23]. Both materials crystallize in the monoclinic structure (the space group $P2_{1}/c$, No. 14) with two formula units in the primitive cell [55, 56]. V4+ has one $d$ electron, while Cu2+ has one $d$ hole. According to the local-density approximation (LDA), the electronic structure of VF4 and CuF2 near the Fermi level is featured by well isolated half-filled band. Therefore, in the first approximation, the properties of these materials can be described by the one-orbital model with the parameters derived from the first-principles calculations. This is basically an extension of the previous considerations to the case of lower crystallographic symmetry. The details of electronic structure calculations and construction of the model are described in Appendix B. The main results are summarized in Table 1.
Table 1: Experimental lattice parameters ($a$, $b$, and $c$ are in Å, $\beta$ is in degrees) and main parameters of the one-orbital model for VF4 and CuF2: the nearest-neighbor hopping $t_{1}$ (in meV), the sign-alternating component of the spin-orbit coupling $t_{y}$ (in meV), and on-site Coulomb repulsion $U$ (in eV). The lattice parameters are taken from Refs. [55] and [56].
$a$ $b$ $c$ $\beta$ $t_{1}$ $t_{y}$ $U$
VF4 $5.381$ $5.170$ $5.340$ $59.74$ 100.66-100.66 $0.71$ $3.2$
CuF2 $3.297$ $4.562$ $4.616$ $83.29$ 173.73-173.73 $5.02$ $4.0$
The primitive translations are ${\bf T}_{1}=(a\sin\beta,0,a\cos\beta)$, ${\bf T}_{2}=(0,b,0)$, and ${\bf T}_{3}=(0,0,c)$, both for VF4 and CuF2. Two sublattices are transformed to each other by the symmetry operation $\{{\cal C}_{2y}|{\bf t}\}$. Therefore, the $x$ and $z$ components of SO interaction between the nearest neighbor have the same sign in all the bonds and can be eliminated (see Appendix B). The $y$ component is sign-alternating. The Néel field parallel to the $y$ axis will result in the weak ferromagnetism along $x$ and $z$.
The atoms in the nearest neighborhood of each atomic site are located at ${\bf t}_{3}$, -${\bf t}_{3}$, ${\bf t}_{3}$-${\bf T}_{2}$, and -${\bf t}_{3}$${\bf T}_{2}$. The AFM propagation vector can be taken as $\boldsymbol{q}=\boldsymbol{G}_{2}$, so that $\sin\frac{\boldsymbol{q}\cdot{\bf t}_{3}}{2}=1$ and $\sin\frac{\boldsymbol{q}\cdot({\bf t}_{3}-{\bf T}_{2})}{2}=-1$. Furthermore, in the first two bonds, the SO coupling is $it_{y}$, while in the second two bonds it is $it_{y}$. Then, using Eqs. (10) and (8) for $\alpha=\frac{\pi}{2}$, and shifting the $\boldsymbol{k}$-mesh $\boldsymbol{k}\to\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$, one can find $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\cos\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$ and $h^{y}_{\boldsymbol{k}}=4t_{y}\sin\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\sin\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$. Finally, $h^{\phantom{0}}_{\boldsymbol{k}}$ in Eq. (16) is the Fourier image of transfer integrals between atoms of one magnetic sublattice and assuming that the they are the same in both sublattices (otherwise the generalized Bloch theorem does not apply). Technically, $h^{\phantom{0}}_{\boldsymbol{k}}$ is obtained by averaging the transfer integrals over two magnetic sublattices. The details and applicability of this approximation and discussed in Appendix B.
Then, using Eq. (28), it is straightforward to find that
$\Omega_{\boldsymbol{k}}^{x}=\frac{Bt_{1}t_{y}}{A_{\boldsymbol{k}}^{3}}(a\cos\beta-c)b\left\{\cos\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})-\cos\boldsymbol{k}\cdot{\bf T}_{2}\right\},$ (40)
$\Omega_{\boldsymbol{k}}^{y}=0$, and
$\Omega_{\boldsymbol{k}}^{z}=-\frac{Bt_{1}t_{y}}{A_{\boldsymbol{k}}^{3}}ab\sin\beta\left\{\cos\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})-\cos\boldsymbol{k}\cdot{\bf T}_{2}\right\}.$ (41)
Therefore, the ratio
$\frac{\Omega_{\boldsymbol{k}}^{x}}{\Omega_{\boldsymbol{k}}^{z}}=\frac{c-a\cos\beta}{a\sin\beta}$ (42)
is controlled by the geometrical factor, depending only on the lattice parameters.
These tendencies are clearly seen in the behavior of AHE and orbital magnetization (see Fig. 5).
Figure 5: Band filling dependence of anomalous Hall conductivity and orbital magnetization in VF4 and CuF2.
The nonvanishing $yz$ and $xy$ components of the Hall conductivity are related to each other by a scaling transformation, which is expected from Eq. (42). The same holds for the $x$ and $z$ components of the orbital magnetization. Furthermore, the shape of orbital magnetization basically repeats the one of AHE, as expected from Eq. (38).
Finally, we would like to comment on the possibility of practical realization of AHE in VF4 and CuF2. One important question, which is typically ignored in prediction of possible altermagnetic materials, is whether $\boldsymbol{q}=\boldsymbol{G}_{2}\equiv(0,1,0)$ corresponds to the true magnetic ground states of VF4 and CuF2. For instance, the analysis of superexchange interactions in the insulating state at the half filling, performed along the same line as for CuO [57], suggests the energy of these interactions has a minimum at $\boldsymbol{q}=(\frac{1}{2},0,\frac{1}{2})$ rather than $\boldsymbol{q}=(0,1,0)$. In the ground state with $\boldsymbol{q}=(\frac{1}{2},0,\frac{1}{2})$, which is expected in VF4 and CuF2, the AHE will vanish and these materials cannot be classified as altermagnets. The same happens in CuO. From the viewpoint of symmetry, CuO could be classified as a potential altermagnet: the space group is $C2/c$, where there are two Cu sublattices. Each sublattice is transformed to itself by the spatial inversion. Different sublattices are transformed to each other by the twofold rotation about $y$, combined by a lattice shift (i.e. similar to VF4 and CuF2). However, the superexchange interactions are such that CuO tends to form a magnetic superstructure with $\boldsymbol{q}\approx(\frac{1}{2},0,\frac{1}{2})$ to become a multiferroic material [57, 58].
VI.3 $\rm RuO_{2}$}-type systems
RuO2 is crystallized in body-centered tetragonal (bct) structure with two sublattices (the space group $P4_{2}/mnm$, No. 136). The atomic positions in both sublattices can be formally generated by the vectors ${\bf t}_{1}=(-\frac{1}{2},\frac{1}{2},\frac{1}{2})$, ${\bf t}_{2}=(\frac{1}{2},-\frac{1}{2},\frac{1}{2})$, and ${\bf t}_{3}=(\frac{1}{2},\frac{1}{2},-\frac{1}{2})$, in units of $a$ and $c$ ($c/a<1$ being the tetragonal distortion along $z$). The additional complication arising in this case is that there are two sign-alternating components of the SO coupling in the nearest bonds: $x$ and $y$, as explained in Fig. 6.
Figure 6: Spin-orbit interactions $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ (denoted by bold blue arrows) around two magnetic sites in the $P4_{2}/mnm$ structure. The vectors $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ replicate the form of Dzyaloshinskii-Moriya interactions. The corresponding bond directions are shown by grey arrows.
The third ($z$) component is identically equal to zero. Thus, formally speaking, there is no weak spin ferromagnetism for the $P4_{2}/mnm$ symmetry.
We assume that the Néel field is parallel to $x$. Then, according to Eqs. (7) and (8), in the local coordinate frame specified by $\boldsymbol{q}=\boldsymbol{G}_{1}$ and $\alpha=0$, the SO interaction is described by the Pauli matrices $\hat{\sigma}_{y}$ and $\hat{\sigma}_{z}$. Other transformations are given by Eqs. (10) and (15).
Therefore, we will deal with the following Hamiltonian:
$\hat{\cal H}_{\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}\hat{\mathbb{1}}+h^{1}_{\boldsymbol{k}}\hat{\sigma}_{x}-h^{y}_{\boldsymbol{k}}\hat{\sigma}_{y}+\left(B+h^{z}_{\boldsymbol{k}}\right)\hat{\sigma}_{z},$ (43)
where it is again convenient to shift the $\boldsymbol{k}$-mesh: $\boldsymbol{k}\to\boldsymbol{k}+(\pi,0,0)$. Then, the parameters of Eq. (43) will be given by $h^{\phantom{0}}_{\boldsymbol{k}}=2(t_{2}-\delta t_{2})(\cos k_{x}+\cos k_{y})+2(t_{2}+\delta t_{2})\cos k_{z}$, $h^{1}_{\boldsymbol{k}}=8t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$, $h^{y}_{\boldsymbol{k}}=-8t_{x}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\sin\frac{k_{z}}{2}$, and $h^{z}_{\boldsymbol{k}}=8t_{y}\cos\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$. Although for the $P4_{2}/mnm$ symmetry $t_{x}=t_{y}$, it is convenient to treat $t_{x}$ and $t_{y}$ as independent parameters. The eigenvalues, eigenfunctions, and properties of Hamiltonian (43) can be obtained along the same line as in Sec. V after replacing $B$ by $B+h^{z}_{\boldsymbol{k}}$.
Furthermore, we have the following properties: $h^{\phantom{0}}_{-\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{-\boldsymbol{k}}=h^{1}_{\boldsymbol{k}}$, $h^{y}_{-\boldsymbol{k}}=h^{y}_{\boldsymbol{k}}$, and $h^{z}_{-\boldsymbol{k}}=-h^{z}_{\boldsymbol{k}}$, which immediately yields: $\varepsilon^{\pm}_{\boldsymbol{k}}(B)=\varepsilon^{\pm}_{-\boldsymbol{k}}(-B)$ and $\varepsilon^{\pm}_{\boldsymbol{k}}(B,t_{y})=\varepsilon^{\pm}_{\boldsymbol{k}}(-B,-t_{y})$. Then, the Berry curvature satisfies the following properties: $\Omega_{\boldsymbol{k}}(B)=-\Omega_{-\boldsymbol{k}}(-B)$ ($\partial_{k_{a}}\phi_{\boldsymbol{k}}$ changes the sign for $a=$ $z$ and $x$, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\theta_{\boldsymbol{k}}$ do not change the sign) and $\Omega_{\boldsymbol{k}}(B,t_{y})=-\Omega_{\boldsymbol{k}}(-B,-t_{y})$ ($\partial_{k_{a}}\theta_{\boldsymbol{k}}$ changes the sign, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\phi_{\boldsymbol{k}}$ do not change the sign). Therefore, the Hall conductivity, $\sigma_{zx}$, will be odd in $B$, $\sigma_{zx}(-B)=-\sigma_{zx}(B)$, and even in $t_{y}$, $\sigma_{zx}(-t_{y})=\sigma_{zx}(t_{y})$. Moreover, $\sigma_{zx}$ is odd in $t_{x}$: $\sigma_{zx}(-t_{x})=-\sigma_{zx}(t_{x})$ (i.e., for the SO component, which is parallel to the Néel field) [26]. Thus, in order to calculate $\sigma_{zx}$ to the first order in the SO coupling, $t_{y}$ can be neglected. This is consistent with the general strategy for calculating DM interactions using the response theory: in order to calculate some particular (say, $x$) component of the DM vectors, it is sufficient to align the exchange field along $x$ and consider only the $x$ component of the SO coupling [30, 59].
After neglecting $t_{y}$, the Berry curvature can be obtained from Eq. (28) as
$\Omega^{y}_{\boldsymbol{k}}=\frac{4t_{1}t_{x}B}{A^{3}_{\boldsymbol{k}}}\cos^{2}\frac{k_{y}}{2}\left(\cos k_{z}-\cos k_{x}\right).$ (44)
The results are summarized in Fig. 7, using parameters reported in Ref. [51].
Figure 7: Results for RuO2-type systems with the parameters $t_{1}=1$, $t_{2}=1.53$, $\delta t_{2}=1.76$, and $t_{x}=0.1$ reported in Ref. [51], and $B/t_{1}=8$: (a) Berry curvature, (b) orbital magnetization for $n_{\rm el}=0.5$ electrons, and (c) band dispersion along high-symmetry directions of bct Brillouin zone (the Fermi level for $n_{\rm el}=0.5$ is shown by dash-dotted line); Corresponding contour plots of (d) Berry curvature and (c) orbital magnetization in the $k_{y}=0$ plane ($0\leq k_{x}\leq 2\pi$ and $0\leq k_{z}\leq 2\pi$); (f) Band filling dependence of anomalous Hall conductivity and orbital magnetization.
The largest contributions to $\Omega^{y}_{\boldsymbol{k}}$ arise from the $(2\pi,0,0)\to(0,0,0)\to(0,0,2\pi)$ directions in the $k_{y}=0$ plane, which are additionally modulated along $k_{y}$, as described by Eq. (44). Moreover, $\Omega^{y}_{\boldsymbol{k}}$ has nodal lines along $k_{z}=\pm k_{x}$. Therefore, if the bands along $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$ were equally populated, $\sigma_{xy}$ would vanish. That is why we need the “tetragonal strain” $\delta t_{2}$ [26], which in the present case is nothing but the tetragonal distortion $c/a<1$ of the bct lattice. It leads to considerably stronger dispersion along $(0,0,0)\to(0,0,2\pi)$. Then, for the $n_{\rm el}=0.5$ electrons, the “majority-spin” band is fully populated along $(0,0,0)\to(2\pi,0,0)$ and only partially populated along $(0,0,0)\to(0,0,2\pi)$, resulting in finite $\sigma_{xy}$. In the expression (38) for ${\cal M}^{y}_{\boldsymbol{k}}$, $\Omega^{y}_{\boldsymbol{k}}$ is additionally modulated by $h^{\phantom{0}}_{\boldsymbol{k}}$, thus destroying the perfect cancellation in between the directions $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$. Unlike in the previous two examples, there is a substantial deviation in the from of the anomalous Hall conductivity and orbital magnetization in Fig. 7(f). This effect is related to the much stronger tetragonal distortion, $\delta t_{2}/t_{1}=1.76$ [51], when the term $h^{\phantom{0}}_{\boldsymbol{k}}\Omega^{y}_{\boldsymbol{k}}$ in the orbital magnetization starts to play a sizable role.
VII Summary and Outlook
The time-reversal symmetry breaking in AFM substances implies certain analogies with the ferromagnetism. On the one hand, it lifts the Kramers degeneracy, splitting the AFM bands with different spin quantum numbers, like in the FM systems. This is an overt analogy, which is intensively explored in the context of altermagnetism [2, 3, 4], but relies on more or less standard picture of antiferromagnetism where there are two magnetic sublattices with opposite directions of magnetic moments.
In the present work, we have argued that there is a deeper analogy, which allows us to present certain classes of centrosymmetric antiferromagnets with broken $\mathcal{T}$ as if they were ferromagnets with only one magnetic site in the unit cell.
Such a presentation appears to be possible due to the hidden $\{\mathcal{S}|{\bf t}\}$ symmetry of the SO interaction in the centrosymmetric antiferromagnets, specifying the transformation to the local coordinate frame where all the spins are pointed along positive direction of $z$, i.e. like in ferromagnets. This symmetry is the consequence of the inversional and translational invariance of the system. The key point is that the inversion symmetry can be broken locally, in individual bonds. Nevertheless, in the centrosymmetric antiferromagnets, $\mathcal{I}$ remains to be one of the symmetry operations of the magnetic lattice. Therefore, if the system is subjected to the lattice distortion, in order to preserve $\mathcal{I}$, this distortion must be antiferroelectric. The antiferroelectric distortion inevitably leads to the doubling of the crystallographic cell, so that the SO interaction behaves as an antiferromagnetically ordered object and changes its sign when it is calculated around the sites of different antiferroelectric sublattices. Nevertheless, the sign change of the SO interaction parameters can be compensated by the similar sign change of the spin Pauli matrices in the local coordinate frame.
Thus, the $\{\mathcal{S}|{\bf t}\}$ symmetry justifies the use of the generalized Bloch theorem, which is widely employed in calculations of incommensurate spin-spiral textures without the SO coupling [28, 49]. Using the generalized Bloch theorem, any such incommensurate texture with the propagation vector $\boldsymbol{q}$ can be described within the crystallographic cell. The AFM alignment of spins is a particular case of the spin spiral, where $\boldsymbol{q}$ is the half of the reciprocal translation. However, the generalized Bloch theorem is typically at odds with the relativistic SO interaction, which tends to lock the spins along some easy axis and severely restrict their ability to rotate on the lattice [59]. The antiferromagnetism appears to be very special case: even though the generalized Bloch theorem is not valid for an arbitrary propagation $\boldsymbol{q}$, it is valid for the AFM propagation, where the antiferroelectric lattice distortion imposes a special symmetry constrain on the SO interaction. As the result, in the local coordinate frame, the SO interaction part responsible for AHE and orbital magnetism, becomes translationally invariant on the “ferromagnetic” lattice with one cite in the unit cell. The same holds for the Néel field.
The regular AFM order doubles the unit cell. In unconventional antiferromagnets with broken $\mathcal{T}$, the magnetic unit cell is expected to coincide with the crystallographic one [1]. We substantially revise this canonical point of view and show that, due to the $\{\mathcal{S}|{\bf t}\}$ symmetry, the magnetic unit cell of centrosymmetric antiferromagnets can be even smaller than the crystallographic one. The situation is highly unusual, but reflects the fundamental symmetry of these systems and naturally explains their similarity with ferromagnets. The emergence of AHE and orbital magnetization in this case becomes natural and they can be evaluated in the same way as for the regular ferromagnets, but in the local coordinate frame, specified by the symmetry operation $\{\mathcal{S}|{\bf t}\}$.
We would like to emphasize that the considered effect has nothing to do with the altermagnetic splitting of bands in centrosymmetric antiferromagnets. From the viewpoint of symmetry, these are different phenomena: the $\{\mathcal{S}|{\bf t}\}$ symmetry is a consequence of the spatial inversional and translational invariance, while the band splitting occurs due to rotational $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry. The latter does not break the $\{\mathcal{S}|{\bf t}\}$ symmetry of the SO interaction itself, which operates between the sites belonging to different antiferroelectric sublattices. The main effect of the $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry on the SO interaction is to fix the signs of different components of this interaction around each magnetic site. However, $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ imposes a symmetry constraint also on the hopping parameters operating within the same sublattices by transforming the hoppings in one magnetic sublattice to “rotated” hoppings in another sublattice. This can split the AFM bands and eventually break the $\{\mathcal{S}|{\bf t}\}$ symmetry.
If the microscopic Hamiltonian were real, the transformations $\{\mathcal{S}|{\bf t}\}$ and $\{\mathcal{T}|{\bf t}\}$ would be identical. However, when the bonds connecting different sublattices are noncentrosymmetric, as in the antiferroelectrically distorted lattice, the Hamiltonian is complex. Then, $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$. Therefore, the time-reversal symmetry can be broken, but the Hamiltonian remain invariant under $\{\mathcal{S}|{\bf t}\}$, making the bands degenerate. This is the fundamental reason why the centrosymmetric antiferromagnets can exhibit AHE and related phenomena. The band splitting provides an additional contribution to AHE. However, it is expected to be relatively small [26].
Thus, the altermagnetic splitting of bands is certainly an interesting theoretical discovery leading to a number of practically important phenomena such as the spin-current generation in AFM substances [15, 18, 24]. However, regarding the whole spectrum of properties expected in centrosymmetric antiferromagnets with broken time-reversal symmetry, it would not be right to attribute all of them to the band splitting and try to consider all of them only from the viewpoint of this splitting. Even if the band splitting is small, as was recently observed in some potential altermagnets [60], the material can still host robust AHE and net orbital magnetization, exhibit magneto-optical rotations and other “ferromagnetic” phenomena, which are related to other, hidden, symmetries of such AFM state, being the consequence if inversional and translational invariance.
Acknowledgement
I am indebted to Sergey Nikolaev and Akihiro Tanaka for collaboration on earlier stages of this project [26]. MANA is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan.
The author declares no competing interests.
Appendix A Elimination of weak ferromagnetic components of the spin-orbit coupling
To be specific, let us consider the case where $x$ and $z$ components of the SO coupling between the sublattices $1$ and $2$ have the same sign in all neighboring bonds, while the $y$ component is sign-alternating, which is rather common for two-sublattice antiferromagnets [26]. The complex transfer integrals, including the SO coupling, are given by $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$.
Then, let us consider the unitary transformation of the sublattice $2$, $\hat{U}=e^{i\varphi\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}}=\cos\frac{\varphi}{2}\hat{\mathbb{1}}+i\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}\sin\frac{\varphi}{2}$, such that $\left(t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}\right)\hat{U}\approx\tilde{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+it^{y}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\sigma}_{y}$. In principle, by using such transformation, one can completely eliminate the SO term $i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ separately in each bond [29, 50]. However, this cannot be done simultaneously for all the bonds, unless the parameters of the SO coupling are the same in all these bonds, as the $x$ and $z$ components of the SO coupling responsible for the weak ferromagnetism, for which $t^{x}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{x}$ and $t^{z}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{z}$.
Thus, assuming $\boldsymbol{n}=\frac{(t_{x},0,t_{z})}{\sqrt{t_{x}^{2}+t_{z}^{2}}}$, which does not depend on the alternating $y$-component, one can find the following equations for $\varphi$:
$t_{x}\cos\frac{\varphi}{2}+n_{x}t\sin\frac{\varphi}{2}-[\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\times\boldsymbol{n}]_{x}\sin\frac{\varphi}{2}=0$ (45)
and
$t_{z}\cos\frac{\varphi}{2}+n_{z}t\sin\frac{\varphi}{2}-[\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\times\boldsymbol{n}]_{z}\sin\frac{\varphi}{2}=0,$ (46)
where $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\times\boldsymbol{n}]_{x}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}^{y}n_{z}$ and $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\times\boldsymbol{n}]_{z}=-t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}^{y}n_{x}$. The first two terms of these equations are of the first order in the SO coupling, while the third terms is of the second order and can be neglected (note that $\varphi$ is of the first order). Therefore, one can find that $\varphi=-2\arctan\frac{\sqrt{t_{x}^{2}+t_{z}^{2}}}{t}$.
Appendix B Electronic structure and minimal model for VF4 and CuF2
The electronic structure of VF4 and CuF2 in LDA with the experimental lattice parameters [55, 56] is shown in Fig. 8.
Figure 8: Electronic structure of VF4 (left) and CuF2 (right) in LDA. The Fermi level is at zero energy. Shaded areas depict the bands, which were used for the construction of the one-orbital models.
The calculations are performed using the linear muffin-tin orbital (LMTO) method [61, 62] and Vosko-Wilk-Nusair parametrization for the exchange-correlation potential in LDA [63].
In both cases, the electronic structure is featured by two well isolated bands at the Fermi level, which are depicted in Fig. 8. For each spin, these bands can accommodate one electron. Therefore, in non-magnetic LDA, the bands are half-filled.
The Bloch functions for these bands can be transformed to the localized Wannier functions, which can be used as the basis states for the minimal model. For the construction of the Wannier functions themselves we use the projector-operator technique [64, 65]. Then, the transfer integrals $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ can be identifies with the matrix elements of LDA Hamiltonian with the SO coupling in the Wannier basis. The screened on-site Coulomb repulsion $U$ can be evaluated within constrained random-phase approximation [66], as explained in Ref. [65]. At the half-filling, the Néel field is related to $U$ as $B\approx U/2$.
The transfer integrals operating within the same sublattice were additionally averaged over two sublattices. Namely, using the transfer integrals in the sublattices $I$ and $II$, one can define $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}+t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the regular contribution) and $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}-t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the altermagnetic contribution). The approximation consists in neglecting $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. This artificially suppress the splitting of the AFM bands, but allows us to employ the generalized Bloch theorem and describe these AFM systems as they would have only one magnetic site in the unit cell. The validity of this approximation depends on the system: while the parameters $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be comparable to $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of CuF2, they appear to be much smaller than $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of VF4 (see Fig. 9).
Figure 9: Distance-dependence of regular ($t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$) and altermagnetic ($\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$) transfer integrals operating within the same sublattices in VF4 and CuF2.